I don't know how to write an equation for this. Please help due today!
How many different license plates are possible if each contains three letters followed by three digits? The letters may repeat although the letters I and O are not used and the digits may repeat?

Found 2 solutions by richwmiller, dabanfield:
Answer by richwmiller(17219)

So why did you wait so long? You had Friday, Saturday and Sunday.

Answer by dabanfield(803)

I don't know how to write an equation for this. Please help due today!
How many different license plates are possible if each contains three letters followed by three digits? The letters may repeat although the letters I and O are not used and the digits may repeat?
We have 26-2 = 24 letters to choose from so there are 24*24*24 ways to select the three letters. There are 10 numbers to choose from so there are 10*10*10 ways to select the three numbers. The number of ways to select three letters followed by three numbers then is:
(24^3)*(10^3) = 24*24*24*10*10*10